# Out-of-distribution generalization via composition: a lens through induction heads in Transformers

### v1

@@ Abstract
Large language models (LLMs) such as GPT-4 sometimes appear to be creative, solving novel tasks often with a few demonstrations in the prompt. These tasks require the models to generalize on distributions different from those from training data—which is known as out-of-distribution (OOD) generalization. Despite the tremendous success of LLMs, how they approach OOD generalization remains an open and underexplored question. We examine OOD generalization in settings where instances are generated according to hidden rules, including in-context learning with symbolic reasoning. Models are required to infer the hidden rules behind input prompts without any fine-tuning.
We empirically examined the training dynamics of Transformers on a synthetic example and conducted extensive experiments on a variety of pretrained LLMs, focusing on a type of components known as induction heads. We found that OOD generalization and composition are tied together—models can learn rules by composing two self-attention layers, thereby achieving OOD generalization. Furthermore, a shared latent subspace in the embedding (or feature) space acts as a bridge for composition by aligning early layers and later layers, which we refer to as the common bridge representation hypothesis.
@@ 1 Introduction
The classical notion of statistical generalization does not seem to account for the progress observed in LLMs. Traditionally, both training instances and test instances are drawn from the same distribution, and it is generally not expected that a model will generalize well on a different test distribution without explicit domain adaptation involving model updates.
The recent success of LLMs suggests a different story: if test data involve compositional structures, LLMs can generalize across different distributions with just a few demonstrations in the prompt (few-shot learning) or even without any demonstrations (zero-shot learning) [16] without updating the model parameters. Indeed, the apparent ability of models to infer rules from the context of a prompt—known as in-context learning (ICL)—is a hallmark of LLMs [21]. Moreover, a growing body of literature on chain-of-thought prompting explicitly exploits the compositional structures of reasoning tasks. For example, phrases like “let’s think step by step” are concatenated with input prompts to elicit reasoning [68, 41], and prompts with intermediate steps are used to improve accuracy on mathematical tasks [86].
@@
  1. 1.
    On the synthetic task of copying sequences of arbitrary patterns, a 2-layer Transformer exhibits an abrupt emergence of subspace matching that accompanies OOD generalization between two Transformer layers, a phenomenon that echoes emergent abilities [86].
  2. 2.
    On language reasoning tasks where LLMs infer the meanings of planted symbols, including examples of in-context learning, OOD generalization requires a similar compositional structure. Extensive experiments on LLMs suggest the presence of a latent subspace for compositions in multilayer and multihead models, which we propose as the common bridge representation hypothesis.
@@ 1.1 An exemplar: copying
We train a 2-layer attention-only Transformer on batches of fresh samples, namely each training step uses independently drawn sequences. Model architecture and training follow the standard practice; see Section B for details. We report both in-distribution (ID) test errors and OOD test errors by computing the average token-wise prediction accuracy based on the second segment. Here the OOD error is evaluated on sequences of a similar format but $\mathcal{P}$ is replaced by the uniform distribution $\mathcal{P}_{\mathrm{ood}}$, and $L$ is replaced by $L_{\mathrm{ood}}=25$. As a comparison, we train a 1-layer attention-only Transformer using the same data.
@@ 2 Dissecting sharp transition in synthetic example
As in Section 1.1, we train a minimally working Transformer for the copying task as a clean synthetic example, though we find similar results on larger models. In particular, the model has 2 self-attention layers with a single head and no MLPs. The architecture and training are standard, including residual connection, LayerNorm, RoPE, dropout, autoregressive training with the AdamW optimizer, and weight decay. See Section A for a list of notations and Section B for experimental details.
Model.
For analytical purposes, we introduce a basic form of Transformer from the circuits perspective [26]. Let ${\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{T}]^{\top}\in\mathbb{R}^{T\times d}$ be the initial embeddings or the hidden states representing a sequence of length $T$ in $d$ dimensions. A multihead self-attention (MSA) is a mapping $\mathrm{MSA}({\bm{X}};{\bm{W}}):\mathbb{R}^{T\times d}\to\mathbb{R}^{T\times d}$ given by
$\mathrm{MSA}({\bm{X}};{\bm{W}}):=\underbrace{\phantom{\big{(}}{\bm{X}}\phantom{\big{)}}}_{\begin{subarray}{c}\text{residual stream stores}\\ \text{info from previous layer}\end{subarray}}+~{}~{}\sum_{j=1}^{H}~{}~{}\overbrace{\mathrm{Softmax}\underbrace{\big{(}{\bm{X}}{\bm{W}}_{\mathrm{QK},j}{\bm{X}}^{\top}\big{)}}_{\begin{subarray}{c}\text{QK circuit reads and}\\ \text{matches info from stream}\end{subarray}}}^{\text{attention matrix}}\underbrace{\phantom{\big{(}}{\bm{X}}{\bm{W}}_{\mathrm{OV},j}^{\top}\phantom{\big{)}}}_{\begin{subarray}{c}\text{OV circuit writes and}\\ \text{adds info to stream}\end{subarray}}$1)
where ${\bm{W}}_{\mathrm{QK},j},{\bm{W}}_{\mathrm{OV},j}\in\mathbb{R}^{d\times d}$ and ${\bm{W}}$ is a collection of all such matrices. An (attention-only) Transformer is a composition of mappings ${\bm{X}}\mapsto\mathrm{MSA}({\bm{X}};{\bm{W}})$ where each layer has different trainable parameters ${\bm{W}}$. Section B.5 provides further details, interpretations, and comparison with practical model variants.
@@ 2.2 Results
Figure 3 shows that a smaller pool size $S$ (decreasing from $1000$ to $750$) requires more training steps to achieve good generalization. Moreover, when $S$ drops below $740$, models fail to generalize well under 20K steps; instead, they may memorize the repetition patterns, yielding small ID errors especially for small $S$. We also find that memorizing models exhibit very different attention matrices (Section D.3), suggesting the failure of learning the copying rule.
@@ 3 Intervention experiments in LLMs
How is the synthetic example relevant to realistic reasoning tasks for LLMs? In this section we address this question by presenting two types of realistic scenarios: out-of-distribution (OOD) prompts (Section 3.1) and realistic compositional structures (Section 3.2). Through these examples, we aim to demonstrate two key points:
@@
We consider a variety of LLMs in our experiments: (1) Llama2-7B [80], (2) Llama3-8B [23], (3) Mistral-7B [38], (4) Falcon-7B [6], (5) Falcon2-11B [50], (6) OlMo-7B [30], (7) Gemma-7B [78], (8) Gemma2-8B [79]. See Appendix C.1 for details about models and implementations.
@@
For each model, we score all attention heads according to Eq. 3 and 4 based on the test prompts, yielding $\mathrm{score}_{\ell,j}^{\mathrm{PTH}}$ and $\mathrm{score}_{\ell,j}^{\mathrm{IH}}$. Then we rank the PTH scores and IH scores in descending order respectively. For a pre-specified $K$, we define PTHs (or IHs) as the top-$K$ attention heads according to $\mathrm{score}_{\ell,j}^{\mathrm{PTH}}$ (or $\mathrm{score}_{\ell,j}^{\mathrm{IH}}$). Section E.2 provides lists of PTHs and IHs in each model.
@@ 3.1 Symbolized language reasoning
In each task, we sample prompts based on a specified rule and use LLMs directly to predict next tokens. Tokens in blue are the target outputs for prediction. See appendix C.2 for a set of complete examples.
@@
    Fuzzy copying: $[A],[B],[C]\ldots[A^{\prime}],[B^{\prime}],{\color[rgb]{0,0,1}[C^{\prime}]}$
@@
    Indirect object identification (IOI): $[\text{Subject}]\ldots[\text{Object}]\ldots[\text{Subject}]\ldots{\color[rgb]{0,0,1}[\text{Object}]}$
@@
    $x_{1},f(x_{1}),x_{2},f(x_{2})\ldots x_{n},{\color[rgb]{0,0,1}f(x_{n})}.$
@@
We measure the prediction accuracy for IOI and ICL in the multiple-choice form, namely we pick the solution with the highest prediction probability among candidate solutions: [Subject] and [Object] for IOI, and the three categories for ICL. Accuracy by random guess is $1/2$ and $1/3$ respectively.
@@
Additionally, on more sophisticated mathematical tasks, we find that chain-of-thought prompting is dependent on IHs (Section E.1).

### v2

@@ Abstract
Large language models (LLMs) such as GPT-4 sometimes appear to be creative, solving novel tasks often with a few demonstrations in the prompt. These tasks require the models to generalize on distributions different from those from training data—which is known as out-of-distribution (OOD) generalization. Despite the tremendous success of LLMs, how they approach OOD generalization remains an open and underexplored question. We examine OOD generalization in settings where instances are generated according to hidden rules, including in-context learning with symbolic reasoning. Models are required to infer the hidden rules behind input prompts without any fine-tuning. We empirically examined the training dynamics of Transformers on a synthetic example and conducted extensive experiments on a variety of pretrained LLMs, focusing on a type of components known as induction heads. We found that OOD generalization and composition are tied together—models can learn rules by composing two self-attention layers, thereby achieving OOD generalization. Furthermore, a shared latent subspace in the embedding (or feature) space acts as a bridge for composition by aligning early layers and later layers, which we refer to as the common bridge representation hypothesis.
@@ 1 Introduction
The classical notion of statistical generalization does not fully explain the progress observed in LLMs. Traditionally, both training instances and test instances are drawn from the same distribution, and it is generally not expected that a model will generalize well on a different test distribution without explicit domain adaptation involving model updates.
The recent success of LLMs suggests a different story: if test data involve compositional structures, LLMs can generalize across different distributions with just a few demonstrations in the prompt (few-shot learning) or even without any demonstrations (zero-shot learning) [16] without updating the model parameters. Indeed, the apparent ability of models to infer rules from the context of a prompt—known as in-context learning (ICL)—is a hallmark of LLMs [20]. Moreover, a growing body of literature on chain-of-thought prompting explicitly exploits the compositional structures of reasoning tasks. For example, phrases like “let’s think step by step” are prepended to input prompts to elicit reasoning [63, 39], and prompts with intermediate steps are used to improve accuracy on mathematical tasks [82].
@@
First, on the synthetic task of copying sequences of arbitrary patterns, a 2-layer Transformer exhibits an abrupt emergence of subspace matching that accompanies OOD generalization between two Transformer layers, a phenomenon that echoes emergent abilities [82].
Second, on language reasoning tasks where LLMs infer the meanings of planted symbols, including examples of in-context learning, OOD generalization requires a similar compositional structure. Extensive experiments on LLMs suggest the presence of a latent subspace for compositions in multilayer and multihead models, which we propose as the common bridge representation hypothesis.
@@ 1.1 An exemplar: copying
$\ldots[A],[B],[C]\ldots[A],[B],{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}[C]}$
@@
We train a 2-layer attention-only Transformer on batches of fresh samples, namely each training step uses independently drawn sequences. Model architecture and training follow the standard practice; see Appendix B for details. We report both in-distribution (ID) test errors and OOD test errors by computing the average token-wise prediction accuracy based on the second segment. Here the OOD error is evaluated on sequences of a similar format but $\mathcal{P}$ is replaced by the uniform distribution $\mathcal{P}_{\mathrm{ood}}$, and $L$ is replaced by $L_{\mathrm{ood}}=25$. As a comparison, we train a 1-layer attention-only Transformer using the same data.
@@ 1.2 Compositional structure is integral to OOD generalization
2 Background
We briefly introduce Transformers, model interpretability, and OOD generalization. For illustrative purposes, we focus on core model modules and selectively present key concepts, deferring further discussion of related work to Section 5 and implementation details to Appendix.
2.1 Transformers
Transformers are standard neural network architectures for processing text and other sequential data. While many variants of Transformers have been developed, their core components align with the original Transformer model proposed by [76]. The first step of processing is tokenization, where a string ${\bm{s}}$ is converted into tokens. Each token is an element of a vocabulary $\mathcal{A}$, typically representing words, subwords, special symbols, and similar units. Then, these tokens are mapped to numerical vectors known as token embeddings. In the original formulation [76], positional information is encoded by a set of positional embedding vectors, and the sum of token and positional embeddings produce input vectors ${\bm{x}}_{1},\ldots,{\bm{x}}_{T}\in\mathbb{R}^{d}$ to a Transformer.
Transformer architecture.
The input vectors are processed by multiple Transformer layers, which yield hidden states ${\bm{x}}_{1}^{(\ell)},\ldots,{\bm{x}}_{T}^{(\ell)}\in\mathbb{R}^{d}$ after each layer $\ell$. Let ${\bm{X}}=[{\bm{x}}_{1}^{(\ell)},\ldots,{\bm{x}}_{T}^{(\ell)}]^{\top}\in\mathbb{R}^{T\times d}$ be the input vectors or the hidden states at a layer $\ell$ ($\ell=0$ means the input vectors). The computations at each layer follow the generic formulas: for each $\ell=1,2,\ldots,L$,
$\begin{split}&{\bm{X}}\longleftarrow{\bm{X}}+\mathrm{MSA}({\bm{X}};{\bm{W}}^{(\ell)})\\ &{\bm{X}}\longleftarrow{\bm{X}}+\mathrm{FFN}({\bm{X}};\widetilde{\bm{W}}^{(\ell)})\end{split}$ (1)
where ${\bm{W}}^{(\ell)}$ and $\widetilde{\bm{W}}^{(\ell)}$ contain all trainable parameters. Here, a crucial component is the multihead self-attention (MSA), which is a mapping $\mathrm{MSA}({\bm{X}};{\bm{W}}):\mathbb{R}^{T\times d}\to\mathbb{R}^{T\times d}$ given by
$\mathrm{MSA}({\bm{X}};{\bm{W}})=\sum_{j=1}^{H}\overbrace{\mathrm{Softmax}\hskip 2.84526pt\big{(}{\bm{X}}{\bm{W}}_{\mathrm{QK},j}{\bm{X}}^{\top}\big{)}}^{\text{attention matrix}}\underbrace{\phantom{\big{(}}{\bm{X}}{\bm{W}}_{\mathrm{OV},j}^{\top}\phantom{\big{)}}}_{\begin{subarray}{c}\text{OV circuit writes and}\\ \text{adds info to stream}\end{subarray}}$2)
where ${\bm{W}}_{\mathrm{QK},j},{\bm{W}}_{\mathrm{OV},j}\in\mathbb{R}^{d\times d}$, and ${\bm{W}}$ denotes the collection of all such matrices. The softmax operation applies to each row of ${\bm{X}}{\bm{W}}_{\mathrm{QK},j}{\bm{X}}^{\top}$ and yields attention weights that are positive and sum to $1$ by definition. The feedforward network $\mathrm{FFN}$ applies to each column of ${\bm{X}}$ individually, and one common form reads $\mathrm{FFN}({\bm{x}}_{t};\widetilde{\bm{W}})=\widetilde{\bm{W}}_{1}\mathrm{ReLU}(\widetilde{\bm{W}}_{2}{\bm{x}}_{t})$ for each position $t$.
At the final layer $L$, there is a simple classification layer $\mathrm{Softmax}({\bm{W}}^{(\mathrm{out})}{\bm{h}}_{T}^{(L)})$ where ${\bm{W}}^{(\mathrm{out})}\in\mathbb{R}^{|\mathcal{A}|\times d}$, to convert the last-position hidden states ${\bm{h}}_{T}^{(L)}$ to a probability vector, which is used to predict the next token. During pretraining, the entire model (weights in all layers and token embeddings) is trained with backpropagation.
Historical development.
Word embedding [50, 49, 59], the precursor to Transformers, aims to represent words as numerical vectors, similar to the input vectors described earlier. To find contextualized representation of text strings, recurrent neural networks (RNNs) and long short-term memory networks (LSTMs) were initially employed, but it was soon found that attention mechanism [77, 10] handles long-range dependencies more effectively, laying the foundation for Transformers. As the scale of model sizes and pretraining corpora increased, Transformers started to exhibit surprising capabilities. Notably, they can generalize on many downstream tasks with crafted prompts without even updating the model parameters [63]. For example, in in-context learning, a string of instructions or examples are added as a prefix to the test string, and the model seemingly “learns” from the context and generalizes on new tasks.
2.2 Interpreting Transformers
How do Transformers process and represent text data? While Transformers are designed and often used as black-box models, recent analyses offer partial insights into this question. The hidden states ${\bm{x}}_{t}$ are believed to encode the semantic meaning of input texts as a linear combination of base concept vectors, e.g., [88]
[…]
which means that the hidden states representing the word “apple” is decomposed into a linear combination of vectors of interpretable base concepts/features. A further refined decomposition of “other” vector may reveal syntactic features, positional features, and other contextual features.
[…]
Below we present a list of key concepts often used for interpreting Transformers.
  • Linear representation hypothesis (LRH) and feature superposition. The LRH posits that concepts are encoded as linear subspaces within the embedding or hidden states space. Feature superposition posits that hidden states are sparse linear combinations of base concept vectors from a potentially large dictionary.
  • Attention matrix. It is the output of $\mathrm{Softmax}$ operation in each of $H$ attention heads in Eqn. 2. An attention matrix ${\bm{A}}\in\mathbb{R}^{T\times T}$ indicates how an attention head aggregates information from past tokens via the weighted sum, where the weights correspond to the entries of ${\bm{A}}$.
  • Residual stream. It is the output of the identity map (namely a copy of ${\bm{X}}$) in Eqn. LABEL:eq:update, which is updated by the outputs of both $\mathrm{MSA}$ and $\mathrm{FFN}$ in each Transformer layer.
  • QK circuit. It is a component of an attention head that computes ${\bm{Z}}={\bm{X}}{\bm{W}}_{\mathrm{QK},j}{\bm{X}}^{\top}\in\mathbb{R}^{T\times T}$. A QK circuit provides a similarity measure between two hidden states, and ${\bm{Z}}$ decides at each token how much an attention head “reads” and “processes” information from past tokens.
  • OV circuit. It is a component of an attention head that computes ${\bm{A}}{\bm{X}}{\bm{W}}_{\mathrm{OV},j}^{\top}$, whose output is added to the residual stream. The OV circuit decides how information is “written” to the stream.
  • Previous-token head (PTH). It refers to a type of attention heads in a Transformer that attends to the previous token, namely $A_{t,t-1}\approx 1$, for a typical input text.
  • Induction head (IH). It refers to a type of attention heads that attends to the token next to an identical token on an input string with repetition patterns.
In the literature, PTHs and IHs are usually qualitative characterization of attention patterns as they are input-dependent. In this paper, we identify PTHs and IHs from all attention heads by calculating and ranking scores based on desired attention patterns, cf. Eqn. 56.
2.3 OOD generalization
Since GPT-3 and GPT-4, many research communities have showed significant interest in understanding why trained models generalize well on novel tasks. Broadly speaking, OOD generalization refers to the generalization behavior of models when training data have a different distribution compared with the test data. Unlike classical statistics topics such as extrapolation, distribution shift, and domain adaptation, pretrained LLMs can solve compositional tasks and seemingly learn from novel context. While mathematical characterization of OOD data for natural languages remains elusive, recent experimental investigations have offered insights. We highlight several representative approaches for generating OOD data.
  1. 1.
    Changing the length or the distribution of task strings. This is particularly useful for experiments with synthetic data, where the distribution of task strings can be easily modified [41, 38, 78].
  2. 2.
    Replacing parts of a text string with counterfactual statements or unnatural symbols. It is used for probing LLMs to test reliability or reveal the internal mechanism [64, 85, 56, 32].
  3. 3.
    Constructing compositional tasks based on deductive rules. It is used for testing models on complex reasoning tasks to examine whether models have learned the latent rules [65].
However, prior studies either fail to probe the inner workings of Transformer models or lack systematic measurements across many models. Compared with existing papers, we provide in-depth measurements to probe model behavior during training dynamics, and comprehensive experiments across many pretrained models, ranging from small Transformer such as GPT-2 (124M parameters) to larger LLMs such as Llama-3 (70B parameters). Moreover, our findings about subspace matching are novel, especially in the context of compositionality of LLMs.
@@ 3 Dissecting sharp transition in synthetic example
As in Section 1.1, we train a minimally working Transformer for the copying task as a clean synthetic example, though we find similar results on larger models. In particular, the model has 2 self-attention layers with a single head and no MLPs. The architecture and training are standard, including residual connection, LayerNorm, RoPE, dropout, autoregressive training with the AdamW optimizer, and weight decay. In the Appendix, we provide a list of notations, experimental details, and comparison with model variants.
@@ 3.2 Results
Figure 3 shows that a smaller pool size $S$ (decreasing from $1000$ to $750$) requires more training steps to achieve good generalization. Moreover, when $S$ drops below $740$, models fail to generalize well under 20K steps; instead, they may memorize the repetition patterns, yielding small ID errors especially for small $S$. We also find that memorizing models exhibit very different attention matrices (Figure 1516), suggesting the failure of learning the copying rule.
@@ 4 Intervention experiments in LLMs
How is the synthetic example relevant to realistic reasoning tasks for LLMs? In this section, we address this question by presenting two types of realistic scenarios: out-of-distribution (OOD) prompts (Section 4.1 and 4.2) and realistic compositional structures (Section 4.3). Through these examples, we aim to demonstrate two key points:
@@
We consider a variety of LLMs in our experiments: (1) Llama2-7B [75], (2) Llama3-8B [22], (3) Mistral-7B [35], (4) Falcon-7B [5], (5) Falcon2-11B [46], (6) OlMo-7B [30], (7) Gemma-7B [72], (8) Gemma2-8B [73]. See Appendix D for details about models and implementations.
@@
For each model, we score all attention heads according to Eq. 5 and 6 based on the test prompts, yielding $\mathrm{score}_{\ell,j}^{\mathrm{PTH}}$ and $\mathrm{score}_{\ell,j}^{\mathrm{IH}}$. Then we rank the PTH scores and IH scores in descending order respectively. For a pre-specified $K$, we define PTHs (or IHs) as the top-$K$ attention heads according to $\mathrm{score}_{\ell,j}^{\mathrm{PTH}}$ (or $\mathrm{score}_{\ell,j}^{\mathrm{IH}}$).
See Appendix D for a set of complete examples we introduce below.
@@ 4.1 Symbolized language reasoning
In each task, we sample prompts based on a specified rule and use LLMs directly to predict next tokens. Tokens in blue are the target outputs for prediction.
@@
    Fuzzy copying: $[A],[B],[C]\ldots[A^{\prime}],[B^{\prime}],{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}[C^{\prime}]}$
@@
    Indirect object identification (IOI):
    $[\text{Subject}]\ldots[\text{Object}]\ldots[\text{Subject}]\ldots{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}[\text{Object}]}$
@@
    $x_{1},f(x_{1}),x_{2},f(x_{2})\ldots x_{n},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}f(x_{n})}.$
@@
Figure 5: Scaling experiment on fuzzy copying: how removal of induction heads impact Pythia models. We use the family of Pythia models of varying sizes, ranging from 36M parameters to 7B parameters. As we increase the number of removed induction heads, there is a consistent accuracy drop of all models on fuzzy copying.
We measure the prediction accuracy for IOI and ICL in the multiple-choice form, namely we pick the solution with the highest prediction probability among candidate solutions: [Subject] and [Object] for IOI, and the three categories for ICL. We report prediction accuracy based on $100$ sampled prompts for each task. Accuracy by random guess is $1/2$ and $1/3$ respectively.
In addition, we examine how the accuracy changes across different scales of models. We use the Pythia model suite [13], which is a collection of Transformers of varying sizes pretrained on the same data.
@@
4.2 Mathematical reasoning with chain-of-thought
GSM8K "Jerry is cutting up wood for his wood-burning stove. Each pine tree makes 80 logs, each maple tree makes 60 logs, and each walnut tree makes 100 logs. If Jerry cuts up 8 pine trees, 3 maple trees, and 4 walnut trees, how many logs does he get?" "First find the total number of pine logs by multiplying the number of trees by the number of logs per tree: 80 logs/pine * 8 pines = <<80*8=640>>640 logs Then do the same thing for the maple trees: 60 logs/maple * 3 maples = <<60*3=180>>180 logs And do the same thing for the walnut trees: 100 logs/walnut * 4 walnuts = <<100*4=400>>400 logs Finally, add up the number of logs from each type of tree to find the total number: 640 logs + 180 logs + 400 logs = <<640+180+400=1220>>1220 logs #### 1220"
GSM8K-rand Template "{name} is cutting up wood for his wood-burning stove. Each {pine} tree makes {pine_logs} logs, each {maple} tree makes {maple_logs} logs, and each {walnut} tree makes {walnut_logs} logs. If {name} cuts up {pine_count} {pine} trees, {maple_count} {maple} trees, and {walnut_count} {walnut} trees, how many logs does he get? "First find the total number of {pine} logs by multiplying the number of trees by the number of logs per tree: {pine_logs} logs/{pine} * {pine_count} {pine} = <<{pine_logs}*{pine_count}={total_pine}>>{total_pine} logs Then do the same thing for the maple trees: {maple_logs} logs/{maple} * {maple_count} {maple} = <<{maple_logs}*{maple_count}={total_maple}>>{total_maple} logs And do the same thing for the walnut trees: {walnut_logs} logs/{walnut} * {walnut_count} {walnut} = <<{walnut_logs}*{walnut_count}={total_walnut}>>{total_walnut} logs Finally, add up the number of logs from each type of tree to find the total number: {total_pine} logs + {total_maple} logs + {total_walnut} logs = <<{total_pine}+{total_maple}+{total_walnut}={total}>>{total} logs #### {total}"
Figure 6: An example from the GSM8K benchmark, and a template from GSM8K-rand that we constructed. Left: a sample question and its solution with CoT reasoning. The final solution is the number $1220$ after ####. Right: we sample variables name, pine, maple, walnut independently from a pool of $40$ symbol strings. Then we sample other variables such as pine_logs and pine_count independently from a range of integers. We ensure the reasoning logic leads to the correct solution total.
To further examine the impact of induction heads in a more realistic setting, we consider a popular benchmark dataset GSM8K [19], which contains a collection of K-12 mathematical questions. Solving the questions requires basic arithmetic abilities—usually a combination of addition, subtraction, multiplication, and division—on top of natural language understanding.
Figure 6 (left) shows an example of a typical question and a solution with chain-of-thought (CoT) reasoning [83]. CoT is a standard technique for LLMs to solve reasoning tasks, where intermediate steps are explicitly shown in in-context examples. Given the prompt that contains these in-context examples, LLMs are observed to be significantly better at solving reasoning tasks. We consider 10-shot learning, where $10$ question-solution pairs are added as a prefix to a test question.
To generate explicit OOD test data, we constructed a variant of GSM8K which we call GSM8K-rand. This dataset comprises $12$ templates derived from maintaining the core structures of $12$ question-solution examples from GSM8K and randomizing names and numbers, as shown in Figure 6 (right). Inspired by [51], we sample the numbers from a range of integers with 1 to 3 digits, and sample the names and items from a predefined pool. A key difference from [51] is that the pool for names and items consists of $40$ tuples of special symbols such as “\$#”, “@@@@%”, “!&”, “*#”, “#@@@@”, so the test examples were unlikely seen by the LLMs during training. Similar to the previous three tasks, we measure the accuracy of models under the removal of induction heads. Appendix D details the template construction and measurements.
Results: CoT reasoning relies heavily on IHs.
Similar to Figure 4, we remove induction heads and measure the accuracy of three LLMs on GSM8K and GSM8K-rand. Table 1 shows that induction heads have a large impact on model accuracy on both datasets, as their removal significantly decrease the model accuracy compared with the removal of the same number of random heads. One possible explanation is that CoT reasoning depends on copying abilities to reuse intermediate outputs in its subsequent calculation. We leave a comprehensive investigation to future work.
Accuracy Number of IHs / random heads removed
0 10 20 30 40 50
Llama2-7B GSM8K 0.11 0.06/0.12 0.01/0.11 0.01/0.09 0.02/0.07 0.01/0.07
GSM8K-rand 0.29 0.23 /0.27 0.10/0.25 0.08/0.19 0.06/0.20 0.08/0.16
0 50 100 150 200 250
Llama2-70B GSM8K 0.59 0.41/0.55 0.21/0.52 0.21/0.52 0.09/0.53 0.01/0.50
GSM8K-rand 0.87 0.68/0.81 0.53/0.80 0.51/0.80 0.29/0.78 0.10/0.76
Llama3-70B GSM8K 0.89 0.88/0.90 0.92/0.90 0.86/0.89 0.60/0.88 0.51/0.87
GSM8K-rand 0.96 0.96/0.96 0.95/0.97 0.61/0.98 0.40/0.97 0.51/0.94
Table 1: CoT reasoning depends on IHs on mathematical tasks. Accuracy comparison under removal of IHs vs. random heads (averaged over 5 random seeds)